# OpenSeg-R: Improving Open-Vocabulary Segmentation via Step-by-Step Visual Reasoning

### v1

@@ 1 Introduction
Despite their success, these methods rely solely on associating class name with visual regions for segmentation, without any interpretable explanation or reasoning for segmentation outcomes. This lack of transparency undermines trust in model decisions and limits adaptability to some complex scenarios and difficult examples. For example, as illustrated in  Fig. 1(a), an open-vocabulary segmentor incorrectly classifies a shelf as a stool due to their similar appearance and semantics.
@@
Concretely, our framework mimics a human-like reasoning process using a Large Multimodal Model (LMM), such as Qwen2.5-VL Qwen2.5-VL . The LMM is prompted to generate Image-Specific Reasoning, which begins by producing an overall description of the image and then identifying the presence of relevant object categories. Based on the observed classes, the LMM progressively explains the association between each object and its corresponding category through coarse-to-fine reasoning steps. As illustrated in Fig. 1(b), the hierarchical reasoning process first identifies the coarse and sub-category levels, which helps eliminate irrelevant distractor classes (e.g., distinguishing food from bulb). Subsequently, fine-grained attribute-based descriptions are used to disambiguate visually similar categories (e.g., shelf vs. stool).
@@ 3.2 Step-by-Step Visual Reasoning
As shown in Fig. 2, our OpenSeg-R first performs step-by-step visual reasoning to generate image-specific reasoning. Next, we apply a Class Aligner to map the class names predicted by the LMM back to the original class set. To complement missing categories, we further introduce Generic Class Reasoning. The combined reasoning outputs are then passed to the Prompt Composer and Segmentor, followed by mask ensembling to produce the final segmentation results. We provide a detailed illustration of each component in the following sections.
@@ 3.2.2 Class Aligner
Although the LMM can recognize most salient objects, it may overlook classes that occupy small regions or have low visual prominence —such as the ‘sidewalk’ in Fig. 2. In addition, when the candidate class list is too long, the LMM may ignore certain valid classes due to its limited context length and attention capacity. Such omissions can result in incomplete or biased segmentation outcomes. To address this issue, we introduce a Generic Class Reasoning module as a complementary component to the image-specific reasoning.
@@ 3.3 Mask Generation
We employ a reason prompt composer ${P}_{r}$ to convert each class-reason pair from $\mathcal{T}$ into a descriptive textual prompt for the segmentor. Unlike traditional prompts that consist only of class names (e.g., ‘a photo of {$c$}’), our ${P}_{r}$ incorporates the associated reasoning to enrich semantic guidance. Specifically, for each pair $(c,r)$, we use a compositional template such as ‘a photo of {$c$} has {$r$}’ to construct the final prompt. Formally, the composed prompts are generated as:
@@ 4.3 Comparison with State-of-the-arts
We compare our method with state-of-the-art approaches on open-vocabulary semantic segmentation, as shown in Tab. 1. Among them, SED xie2024sed and MAFT+ jiao2024collaborative are representative one-stage and two-stage segmentation frameworks, respectively. We integrate our reasoning module into both architectures, resulting in two variants: OpenSeg-R w/ SED and OpenSeg-R w/ MAFT+. With a relatively small vision-language model (VLM), our method consistently outperforms the corresponding baselines across all benchmarks. The proposed Open-OpenSeg-R variants achieve the best results in all settings. Specifically, on A-150, our method yields an improvement of approximately 2% when integrated with SED xie2024sed . On A-847, which contains more fine-grained categories than A-150, OpenSeg-R w/ MAFT+ achieves a 1.7% gain over MAFT+ jiao2024collaborative , surpassing or matching the performance of methods that utilize significantly larger VLMs. A similar trend is observed on PAS-20, further demonstrating the effectiveness and generalization ability of our reasoning-enhanced framework. Our method surpasses all SOTA methods on A-150, A-847, and PC-459.
@@
In Tab. 2, we compare our method on the open-vocabulary panoptic segmentation task. Once again, our approach achieves consistent improvements across all evaluation metrics and outperforms all baselines. Specifically, it brings a gain of 0.9% in Panoptic Quality (PQ), 2.3% in Segmentation Quality (SQ), and 1.3% in Recognition Quality (RQ). These results demonstrate the effectiveness of our reasoning-guided framework in enhancing both semantic understanding and instance-level discrimination in open-vocabulary scenarios.
@@ 4.4 Ablation Study
The reasoning output consists of three components: broad class, sub-class, and fine-grained visual attributes. Based on these, we construct four types of prompts: (1) class name only (baseline), (2) coarse reason (broad + sub-class), (3) coarse + attribute, and (4) attribute only. As shown in Tab. 3, using coarse reasoning alone slightly degrades performance, as such descriptions (e.g., ‘a dog that is an animal’) are too generic to distinguish similar categories. In contrast, incorporating fine-grained attributes consistently improves results across all datasets. Notably, prompts using attributes alone achieve the best performance in most cases, suggesting that detailed visual cues provide the most discriminative and informative guidance for open-vocabulary segmentation.
@@
We evaluate the effectiveness of different reasoning types in Tab. 4, comparing three configurations: (1) generic class reasoning only, (2) image-specific reasoning only, and (3) their combination. Generic reasoning performs worse across most datasets, as its descriptions are often too broad to distinguish visually similar objects, leading to lower segmentation accuracy. In contrast, image-specific reasoning, which is inferred from image content, provides more discriminative cues and achieves better performance. However, LMMs may still overlook certain objects. By combining both reasoning types, our method compensates for such omissions: generic reasoning recovers missed categories, while image-specific reasoning improves class-level discrimination. As shown in the table, this hybrid approach yields the best overall results. For PAS-20, which contains relatively few and easily recognizable categories, image-specific reasoning alone is sufficient, as the LMM can reliably detect all relevant classes without generic supplementation.
@@ 4.5 Qualitative Results
To further illustrate the effectiveness of our method, we present qualitative comparisons in Fig. 3. We present several cases where baseline methods MAFT+ jiao2024collaborative fail to correctly segment or classify objects due to ambiguous visual or semantic cues. For each case, we show the visual reasoning attributes inferred by the LMM. These reasoning cues help disambiguate visually similar categories and guide the segmentor toward more accurate predictions. As shown, our method produces more semantically aligned masks across a variety of categories, such as floor, tree, road, and sand, demonstrating the benefit of our OpenSeg-R in open-vocabulary segmentation.
@@ 5 Conclusion
In this paper, we introduces a novel method OpenSeg-R that improves open-vocabulary segmentation through step-by-step visual reasoning. OpenSeg-R leverages Large Multimodal Models (LMMs) to perform hierarchical reasoning based on the input image, generating both image-specific and generic class explanations, which are then composed into descriptive prompts to guide segmentation. Experiments on open-vocabulary semantic and panoptic segmentation benchmarks demonstrate the effectiveness and broad applicability of our approach.
Despite its promising results, OpenSeg-R has several limitations. The quality of generated reasoning depends on the reliability of the LMM, which may produce ambiguous outputs in complex scenes. In addition, incorporating an LMM increases model complexity and computational overhead. Nevertheless, our framework opens up promising directions for future research. With step-by-step visual reasoning, our method enhances model robustness by effectively reducing ambiguity in visually or semantically similar cases. This design not only improves open-vocabulary segmentation performance but also provides human-understandable explanations, which is an important step toward trustworthy deployment in safety-critical applications.
@@ Appendix A Qualitative Results of Open-Vocabulary Panoptic Sementation
In Fig. 4, we present qualitative comparisons with MAFT+ jiao2024collaborative on open-vocabulary panoptic segmentation. Our method consistently delivers superior performance across most cases. Notably, it demonstrates improved instance-level discrimination. For example, it more accurately identifies ambiguous categories such as stool, and small objects like plates.
@@ Appendix B More Implementation Details
We adopt Qwen2.5-VL-72B-Instruct Qwen2.5-VL as our Large Multimodal Model (LMM) to generate image-specific reasoning. This model is chosen for its strong instruction-following ability and robust image understanding. Its responses are generally well-structured and coherent, which makes it easier to perform downstream processing and reasoning alignment. We provide the detailed prompts used with Qwen2.5-VL-72B-Instruct Qwen2.5-VL to perform step-by-step visual reasoning in our framework. As illustrated in Fig. 5, the reasoning process consists of two components: image-specific reasoning and generic class reasoning. For image-specific reasoning, we guide the LMM through three stages: (1) generating a global image description, (2) identifying candidate classes present in the scene, and (3) producing hierarchical, fine-to-grained reasoning for each observed class. For generic class reasoning—used to supplement missed classes—we similarly apply a three-step process that involves coarse category identification, sub-category refinement, and attribute-based justification. This design ensures that the reasoning output is structured, interpretable, and directly usable for composing descriptive prompts in the segmentation stage. During the use of Qwen2.5-VL-72B-Instruct Qwen2.5-VL , we set the decoding temperature to 0.7. In a few cases, Qwen2.5-VL-72B-Instruct Qwen2.5-VL fails to recognize certain objects in the image despite multiple attempts. For these instances, we discard the image-specific reasoning and rely solely on the generic class reasoning to supplement the segmentation process.

### v2

@@ 1 Introduction
Despite their success, these methods rely solely on associating class name with visual regions for segmentation, without any interpretable explanation or reasoning for segmentation outcomes. This lack of transparency undermines trust in model decisions and limits adaptability to some complex scenarios and difficult examples. For example, as illustrated in  Fig.˜1(a), an open-vocabulary segmentor incorrectly classifies a shelf as a stool due to their similar appearance and semantics.
@@
Concretely, our framework mimics a human-like reasoning process using a Large Multimodal Model (LMM), such as Qwen2.5-VL Qwen2.5-VL . The LMM is prompted to generate Image-Specific Reasoning, which begins by producing an overall description of the image and then identifying the presence of relevant object categories. Based on the observed classes, the LMM progressively explains the association between each object and its corresponding category through coarse-to-fine reasoning steps. As illustrated in Fig.˜1(b), the hierarchical reasoning process first identifies the coarse and sub-category levels, which helps eliminate irrelevant distractor classes (e.g., distinguishing food from bulb). Subsequently, fine-grained attribute-based descriptions are used to disambiguate visually similar categories (e.g., shelf vs. stool).
@@ 3.2 Step-by-Step Visual Reasoning
As shown in Fig.˜2, our OpenSeg-R first performs step-by-step visual reasoning to generate image-specific reasoning. Next, we apply a Class Aligner to map the class names predicted by the LMM back to the original class set. To complement missing categories, we further introduce Generic Class Reasoning. The combined reasoning outputs are then passed to the Prompt Composer and Segmentor, followed by mask ensembling to produce the final segmentation results. We provide a detailed illustration of each component in the following sections.
@@ 3.2.2 Class Aligner
Although the LMM can recognize most salient objects, it may overlook classes that occupy small regions or have low visual prominence —such as the ‘sidewalk’ in Fig.˜2. In addition, when the candidate class list is too long, the LMM may ignore certain valid classes due to its limited context length and attention capacity. Such omissions can result in incomplete or biased segmentation outcomes. To address this issue, we introduce a Generic Class Reasoning module as a complementary component to the image-specific reasoning.
@@ 3.3 Mask Generation
[…]
@@ 4.3 Comparison with State-of-the-arts
We compare our method with state-of-the-art approaches on open-vocabulary semantic segmentation, as shown in Tab.˜1. Among them, SED xie2024sed and MAFT+ jiao2024collaborative are representative one-stage and two-stage segmentation frameworks, respectively. We integrate our reasoning module into both architectures, resulting in two variants: OpenSeg-R w/ SED and OpenSeg-R w/ MAFT+. With a relatively small vision-language model (VLM), our method consistently outperforms the corresponding baselines across all benchmarks. The proposed Open-OpenSeg-R variants achieve the best results in all settings. Specifically, on A-150, our method yields an improvement of approximately 2% when integrated with SED xie2024sed . On A-847, which contains more fine-grained categories than A-150, OpenSeg-R w/ MAFT+ achieves a 1.7% gain over MAFT+ jiao2024collaborative , surpassing or matching the performance of methods that utilize significantly larger VLMs. A similar trend is observed on PAS-20, further demonstrating the effectiveness and generalization ability of our reasoning-enhanced framework. Our method surpasses all SOTA methods on A-150, A-847, and PC-459.
@@
In Tab.˜2, we compare our method on the open-vocabulary panoptic segmentation task. Once again, our approach achieves consistent improvements across all evaluation metrics and outperforms all baselines. Specifically, it brings a gain of 0.9% in Panoptic Quality (PQ), 2.3% in Segmentation Quality (SQ), and 1.3% in Recognition Quality (RQ). These results demonstrate the effectiveness of our reasoning-guided framework in enhancing both semantic understanding and instance-level discrimination in open-vocabulary scenarios.
@@ 4.4 Ablation Study
The reasoning output consists of three components: broad class, sub-class, and fine-grained visual attributes. Based on these, we construct four types of prompts: (1) class name only (baseline), (2) coarse reason (broad + sub-class), (3) coarse + attribute, and (4) attribute only. As shown in Tab.˜3, using coarse reasoning alone slightly degrades performance, as such descriptions (e.g., ‘a dog that is an animal’) are too generic to distinguish similar categories. In contrast, incorporating fine-grained attributes consistently improves results across all datasets. Notably, prompts using attributes alone achieve the best performance in most cases, suggesting that detailed visual cues provide the most discriminative and informative guidance for open-vocabulary segmentation.
@@
We evaluate the effectiveness of different reasoning types in Tab.˜4, comparing three configurations: (1) generic class reasoning only, (2) image-specific reasoning only, and (3) their combination. Generic reasoning performs worse across most datasets, as its descriptions are often too broad to distinguish visually similar objects, leading to lower segmentation accuracy. In contrast, image-specific reasoning, which is inferred from image content, provides more discriminative cues and achieves better performance. However, LMMs may still overlook certain objects. By combining both reasoning types, our method compensates for such omissions: generic reasoning recovers missed categories, while image-specific reasoning improves class-level discrimination. As shown in the table, this hybrid approach yields the best overall results. For PAS-20, which contains relatively few and easily recognizable categories, image-specific reasoning alone is sufficient, as the LMM can reliably detect all relevant classes without generic supplementation.
@@ 4.5 Qualitative Results
To further illustrate the effectiveness of our method, we present qualitative comparisons in Fig.˜3. We present several cases where baseline methods MAFT+ jiao2024collaborative fail to correctly segment or classify objects due to ambiguous visual or semantic cues. For each case, we show the visual reasoning attributes inferred by the LMM. These reasoning cues help disambiguate visually similar categories and guide the segmentor toward more accurate predictions. As shown, our method produces more semantically aligned masks across a variety of categories, such as floor, tree, road, and sand, demonstrating the benefit of our OpenSeg-R in open-vocabulary segmentation.
@@ 5 Conclusion
In this paper, we introduce a novel method OpenSeg-R that improves open-vocabulary segmentation through step-by-step visual reasoning. OpenSeg-R leverages Large Multimodal Models (LMMs) to perform hierarchical reasoning based on the input image, generating both image-specific and generic class explanations, which are then composed into descriptive prompts to guide segmentation. Experiments on open-vocabulary semantic and panoptic segmentation benchmarks demonstrate the effectiveness and broad applicability of our approach.
@@ Appendix A Qualitative Results of Open-Vocabulary Panoptic Sementation
In Fig.˜4, we present qualitative comparisons with MAFT+ jiao2024collaborative on open-vocabulary panoptic segmentation. Our method consistently delivers superior performance across most cases. Notably, it demonstrates improved instance-level discrimination. For example, it more accurately identifies ambiguous categories such as stool, and small objects like plates.
@@ Appendix B More Implementation Details
We adopt Qwen2.5-VL-72B-Instruct Qwen2.5-VL as our Large Multimodal Model (LMM) to generate image-specific reasoning. This model is chosen for its strong instruction-following ability and robust image understanding. Its responses are generally well-structured and coherent, which makes it easier to perform downstream processing and reasoning alignment. We provide the detailed prompts used with Qwen2.5-VL-72B-Instruct Qwen2.5-VL to perform step-by-step visual reasoning in our framework. As illustrated in Fig.˜5, the reasoning process consists of two components: image-specific reasoning and generic class reasoning. For image-specific reasoning, we guide the LMM through three stages: (1) generating a global image description, (2) identifying candidate classes present in the scene, and (3) producing hierarchical, fine-to-grained reasoning for each observed class. For generic class reasoning—used to supplement missed classes—we similarly apply a three-step process that involves coarse category identification, sub-category refinement, and attribute-based justification. This design ensures that the reasoning output is structured, interpretable, and directly usable for composing descriptive prompts in the segmentation stage. During the use of Qwen2.5-VL-72B-Instruct Qwen2.5-VL , we set the decoding temperature to 0.7. In a few cases, Qwen2.5-VL-72B-Instruct Qwen2.5-VL fails to recognize certain objects in the image despite multiple attempts. For these instances, we discard the image-specific reasoning and rely solely on the generic class reasoning to supplement the segmentation process.